# finantial support acknowledgement

### v2

@@ Abstract
Federated Learning, designed to address privacy concerns in learning models, introduces a new distributed paradigm that safeguards data privacy but differentiates the attack surface due to the server’s inaccessibility to local datasets and the change in protection objective–parameters’ integrity. Existing approaches, including robust aggregation algorithms, fail to effectively filter out malicious clients, especially those with non-Independently and Identically Distributed data. Furthermore, these approaches often tackle non-IID data and poisoning attacks separately. To address both challenges simultaneously, we present FedCC, a simple yet novel algorithm. It leverages the Centered Kernel Alignment similarity of Penultimate Layer Representations for clustering, allowing it to identify and filter out malicious clients by selectively averaging chosen parameters, even in non-IID data settings. Our extensive experiments demonstrate the effectiveness of FedCC in mitigating untargeted model poisoning and backdoor attacks. FedCC reduces the attack confidence to a consistent zero compared to existing outlier detection-based and first-order statistics-based methods. Specifically, it significantly minimizes the average degradation of global performance by 65.5%. We believe that this new perspective of assessing learning models makes it a valuable contribution to the field of FL model security and privacy. The code will be made available upon paper acceptance.
Index Terms:
@@
Backdoor Attack, Defense, Federated Learning, Model Poisoning Attack, Non-IID, Robust Aggregation
[…]
=-21pt
I Introduction
Federated Learning (FL) [17] has emerged as a promising privacy-preserving model training approach in response to growing concerns about privacy breaches and data leakage caused by centralized learning. While Machine Learning (ML) and Deep Learning (DL) algorithms are popular for their capability, such as personalized recommendations and accurate forecasts, their centralized nature risks data privacy. FL allows data to remain on client devices but exchanges their local model parameters only, avoiding the need for central data collection.
FedAvg [17] is the first algorithm run on the server to aggregate clients’ model parameters. It involves a server initializing and advertising a global model to all clients, who then train the model with their private data and upload the computed parameters to the server. The server averages these parameters to update the global model. These steps are repeated until a stopping criterion is met. FL is especially suitable for data-sensitive environments since raw data remains on clients, reducing the risk of data leakage on the server or in-between communication.
@@ I Introduction
However, due to its distributed nature, FL is susceptible to model poisoning attacks [12]. The server cannot directly examine local datasets’ data quality or model parameter integrity, allowing compromised clients or attackers to manipulate local models. The attacks can degrade global model performance indiscriminately (untargeted attacks) [8, 2, 18] or cause incorrect predictions on specific inputs (targeted attacks) [21, 23, 3, 25]. Backdoor attacks, a type of targeted attack, are stealthier by maintaining overall performance while misclassifying specific inputs [1].
@@
Some preliminary defenses propose robust aggregation methods, but they may compromise privacy by sharing raw data or exposing data distribution to unreliable parties [6, 22]. Existing defenses often target IID data settings, which are rare in practice. Non-IID data, with variations in class distribution, features, or labels across clients [12], can gradually degrade model performance and complicate the detection of malicious clients [28, 27], as the impact of attacks increases with the degree of non-IID [16, 18].
In this study, we present FedCC, a novel defense mechanism for FL that mitigates untargeted model poisoning and targeted backdoor attacks, even with non-IID data. FedCC leverages the Centered Kernel Alignment (CKA) of Penultimate Layer Representations (PLRs) to distinguish between malicious and benign clients while maintaining data privacy by avoiding sharing raw data or additional information between the server and clients. PLRs contain highly distinguishable features, maximizing the differences in CKA scores between benign and malicious clients, even with non-IID data and no server knowledge of local data.
@@
Empirical experiments show that FedCC outperforms existing defenses that rely on first-order statistics. It significantly improves global accuracy against untargeted attacks (65.5% compared to similar approaches) and reduces attack confidence in targeted backdoor attacks to almost zero while preserving main task accuracy. These results highlight FedCC’s effectiveness and robustness in mitigating attacks in FL settings, particularly in non-IID data settings.
@@
  • We propose FedCC, a simple yet novel detection and aggregation method for FL, to defend against model poisoning attacks (untargeted model poisoning and targeted backdoor attacks) by comparing the CKA of PLRs of local and global models and averaging the selected local model parameters.
@@
  • We justify that PLR is the most distinguishable and indicative feature to classify malicious and benign clients by measuring the distance between two clusters of clients.
@@
  • We empirically show that CKA is the most sensitive and accurate metric to measure similarities between differently trained models, especially when data are non-IID, and the server does not have any knowledge about clients or any centralized data.
@@
  • We empirically demonstrated the effectiveness of FedCC for defending against model poisoning attacks. Our experimental results show that comparing the CKA of PLRs is more reliable in filtering malicious clients out than using first-order statistics. FedCC outperforms existing defenses even when data are non-IID.
II Backgrounds and Related Works
II-A Poisoning Attacks in FL
@@
FL is vulnerable to poisoning attacks due to its distributed nature and the server’s inability to examine dataset quality and model parameter integrity directly. Model poisoning is particularly destructive among various poisoning attacks, as adversaries stealthily manipulate local model parameters to maximize overall performance damage [3, 5]. Thus, this paper focuses on model poisoning attacks.
Model poisoning attacks allow client-side adversaries to alter local model parameters before sending them to the server. To avoid the global model from divergence, attackers optimize the compromised local model for both training loss and an adversarial objective, enhancing its stealthiness [5, 8]. They can also adjust model hyperparameters, such as the learning rate, local epochs, batch size, and optimization objective, to manipulate training rules before and during local training.
Based on the attacker’s goals, model poisoning attacks can be classified as untargeted or targeted. Untargeted attacks aim to degrade overall global test accuracy by manipulating compromised local models. In contrast, targeted attacks aim to mislead the model into misclassifying specific inputs, leaving other classes unaffected. This paper introduces a backdoor ability into targeted model poisoning attacks, where inputs belonging to a chosen class are intentionally misclassified to a chosen label, whereas the performance of other classes remains unaffected. Its added stealthiness makes the attack more challenging to detect [23, 21].
II-B Robust Aggregation in FL
@@
Several Byzantine-robust aggregations based on summary statistics or anomaly detection have been proposed to mitigate model poisoning and backdoor attacks in FL. For example, Krum [4] selected the update with the smallest Euclidean distance to the remaining updates as the new global model. Still, it ignored outliers and assumed IID data only [2]. Median [26] selects the coordinate-wise median of updates, also ignoring outliers. Trimmed Mean [26] excludes extreme values before averaging, and Bulyan [11] combines Krum and Trimmed Mean for selection. These methods rely on knowing the number of attackers, which is often unavailable in practice.
Inspired by the fact that Sybil’s crafted model parameter’s directions are more similar to each other than to benign ones, Foolsgold [9] identified Sybils by measuring cosine similarity between local model gradients, marking the most similar ones as Sybils. However, it struggled with a single malicious client when there were no other malicious clients for comparison. Additionally, this approach may overfit non-IID data, ironically leading to suboptimal performance with IID data. FLTrust [6] used trust scores based on ReLU-clipped cosine similarities to weigh local clients’ parameters differently but risked privacy by requiring clients to share raw data with the server. Furthermore, precise weight scaling is paramount to ensure model convergence.
Previous works often used summary statistics (mean or median), direction (cosine similarity), or distance (Euclidean) of weight vectors, but they struggled to effectively distinguish between benign clients with non-IID data and malicious clients, often misclassifying benign clients as malicious. The lack of performance might be partially attributed to the inadequacy of Euclidean geometry as a metric to compare neural network representations. In contrast, our proposed approach utilizes kernel-based metric (Kernel CKA) to measure the similarity between global and local models, reliably identifying compromised model parameters among the mixture of clients and enhancing defense against model poisoning attacks in FL.
II-C Penultimate Layer Representation in FL
Figure 1: CNN architecture and Penultimate Layer
@@ II-C Penultimate Layer Representation in FL
The penultimate layer is the second-to-last layer (before the softmax layer) of a neural network model, as illustrated in Figure 1. Wang et al. [22] discovered that PLRs are effective in distinguishing malicious models from benign ones; within benign clients, benign PLRs follow the same distribution, whereas malicious PLRs do not, across multiple datasets and neural network architectures. Specifically, they demonstrated that distances among benign PLRs are smaller than those of benign and malicious PLRs. Their proposed method, FLARE, assigned a trust score to each client based on pairwise PLR discrepancies defined by Maximum Mean Discrepancy (MMD) among all model updates, allocating lower values to those farther from the benign distribution. The model updates are then scaled and averaged, weighted by each client’s trust score. However, FLARE continuously redirects the global model using server-owned raw data, increasing the risk of data leakage and jeopardizing data privacy.
II-D Centered Kernel Alignment
Centered Kernel Alignment (CKA) [13] is a highly accurate similarity metric that measures how similar two differently initialized or trained neural networks are, producing a value between 0 (not similar) and 1 (identical). CKA can be calculated by:
$CKA(X,Y)$ $=\frac{HSIC(X,Y)}{\sqrt{HSIC(X,X)HSIC(Y,Y)}}$
$=\frac{{\lVert K(X^{T})K(Y)\rVert}^{2}}{{\lVert K(X^{T})K(X)\rVert}{\lVert K(Y^{T})K(Y)\rVert}}$
@@ II-D Centered Kernel Alignment
where $HSIC$ is Hilbert-Schmidt Independence Criterion [10] (i.e., a not normalized version of CKA) and $K$ is the RBF kernel. It constructs similarity kernel matrices in the weight space and compares these matrices to characterize the representation space, allowing comparisons between layers with varying widths and different initialization schemes.
@@
CKA includes several important properties for similarity measures of neural networks. It is non-invariant to invertible linear transformations, meaning similarity scores change if such transformations are applied. This property is vital; the gradient descent algorithm is not invariant to these transformations because similarity metrics invariant to linear transformations are inaccurate on models trained with gradient descent. It is also invariant to orthogonal transformations and isotropic scaling, which are desirable for neural networks trained with gradient descent owing to their stochastic nature. These properties make CKA a precise and reliable metric for measuring similarities between neural networks, particularly when comparing potential models from malicious clients.
@@
III Threat Model
III-A Attacker’s Goal
We consider two primary attack strategies in the context of FL: untargeted model poisoning and targeted backdoor attacks. Untargeted attacks (Fang attacks [8]) aimed to evade robust aggregation rules like Krum and Coordinate-wise Median or Trimmed Mean. Attackers manipulate the model parameters to degrade the global model’s overall accuracy. Targeted backdoor attacks, represented by Bhagoji et al. [3], aim to deceive the global model into misclassifying specific data samples the attacker chose. The objective is to assign a different label chosen by the attacker while maintaining high accuracy for the remaining classes, making the attack more inconspicuous.
III-B Attackers’ Capabilities
@@
  • An attacker has knowledge about compromised clients’ data samples, current local model parameters, global model parameters of the previous round, and hyperparameters, such as learning rate, optimization method, loss function, batch size, and the number of local epochs.
@@
III-C Attack Strategy
III-C1 Untargeted Model Poisoning Attacks
@@
These attacks based on [8] specifically aim to break Krum and Coordinate-wise Median (Coomed) aggregation rules, known to be byzantine failure tolerant. Their primary objective is to disrupt the model training process, thereby diminishing global test accuracy. The attacks involve manipulating model parameters, such as flipping the sign of malicious parameters, to steer the model in the opposite direction from its uninterrupted trajectory. Specifically, we denote the attack against Krum as Untargeted-Krum and the attack against Coomed as Untargeted-Med.
Untargeted-Krum [8] alters malicious parameters resembling benign ones to maximize the chances of being selected by Krum. Specifically, the optimization problem is as follows:
$\max_{\lambda}\lambda$
subject to $\text{subject to }w_{1}$ $=Krum(w_{1},...,w_{m},w_{m+1},...,w_{n})$
$w_{1}$ $=w_{G}-\lambda s$
$w_{i}$ $=w_{1},\text{ for }i=2,3,...,m$
where $m$ is the number of attackers, $n$ is the total number of selected clients, $w_{G}$ is the global model from the previous iteration, and $s$ is the sign of the global model parameter with no attack. The upper bound of the $\lambda$ is solved as follows:
$\lambda\leq$ $\frac{1}{(n-2m-1)\sqrt{d}}\cdot\min_{m+1\leq i\leq n}\sum_{l\in\Gamma_{w_{i}}^{n-m-2}}D(w_{l},w_{i})$
$+\frac{1}{\sqrt{d}}\cdot\max_{m+1\leq i\leq n}D(w_{i},w_{G})$
@@
where $d$ is the number of parameters in the global model, $\Gamma$ is the set of $n-m-2$ benign local models having the smallest Euclidean distance to $w_{i}$, and $D$ is the Euclidean distance. $\lambda$ is halved until one of the compromised models is selected, or $\lambda$ is less than a threshold.
@@
Untargeted-Med [8] manipulates the model parameters based on their maximum and minimum so that chosen coordinate-wise median values direct toward an inverse direction. The attack starts with defining $w_{max,j}$ and $w_{min,j}$, which indicates the maximum and minimum of the $j$th local model parameters on the benign clients, respectively.
$w_{max,j}$ $=\text{max}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$
$w_{min,j}$ $=\text{min}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$
Also, to avoid sampled $m$ numbers being outliers, if $s_{j}=-1$, $m$ numbers in [$w_{max,j},b\cdot w_{max,j}$] (when $w_{max,j}>0$) or [$w_{max,j},b/w_{max,j}$] (when $w_{max,j}\leq 0$) are randomly sampled. Otherwise, $m$ numbers [$w_{min,j}/b,w_{min,j}$] (when $w_{min,j}>0$) or [$b\cdot w_{min,j},w_{min,j}$] (when $w_{min,j}\leq 0$) are randomly sampled. Since the attack does not depend on $b$, we set $b=2$ as the same as [8].
III-C2 Targeted Backdoor Attack

### v3

# FedCC: Robust Federated Learning against Model Poisoning Attacks

@@ Abstract
Federated learning is a distributed framework designed to address privacy concerns. However, it introduces new attack surfaces, which are especially prone when data is non-Independently and Identically Distributed. Existing approaches fail to effectively mitigate the malicious influence in this setting; previous approaches often tackle non-IID data and poisoning attacks separately. To address both challenges simultaneously, we present FedCC, a simple yet effective novel defense algorithm against model poisoning attacks. It leverages the Centered Kernel Alignment similarity of Penultimate Layer Representations for clustering, allowing the identification and filtration of malicious clients, even in non-IID data settings. The penultimate layer representations are meaningful since the later layers are more sensitive to local data distributions, which allows better detection of malicious clients. The sophisticated utilization of layer-wise Centered Kernel Alignment similarity allows attack mitigation while leveraging useful knowledge obtained. Our extensive experiments demonstrate the effectiveness of FedCC in mitigating both untargeted model poisoning and targeted backdoor attacks. Compared to existing outlier detection-based and first-order statistics-based methods, FedCC consistently reduces attack confidence to zero. Specifically, it significantly minimizes the average degradation of global performance by 65.5%. We believe that this new perspective on aggregation makes it a valuable contribution to the field of FL model security and privacy. The code will be made available upon acceptance.
Keywords:
@@
Federated learning model poisoning attack backdoor attack robust aggregation
1 Introduction
Federated Learning (FL) [22] is a distributed model training framework designed to preserve privacy. FL allows data to remain on client devices but exchanges their local model parameters only for model training. It thus reduces (data) privacy risks associated with centralized learning, not requiring central data collection. Thus, FL is especially suitable for data-sensitive environments since raw data remains on clients, reducing the risk of data leakage on the server or in between communication.
@@ 1 Introduction
However, its distributed nature makes FL susceptible to model poisoning attacks [12]. The server cannot directly examine local datasets’ data quality or model parameter integrity, leaving compromised clients or attackers to manipulate local models. This can degrade global model performance indiscriminately (untargeted attacks) [7, 2, 26] or cause incorrect predictions on specific inputs (targeted attacks) [29, 31, 3, 33]. Backdoor attacks are stealthier targeted attacks that maintain overall performance while misclassifying specific inputs [1].
@@
Some preliminary defenses propose robust aggregation methods, but they may compromise privacy by sharing raw data or exposing data distribution to untrusted parties [6, 30]. While some defenses exist, most focus on IID data settings, which are rare in real-world scenarios. Non-IID data, with variations in class distribution and features across clients [12], can gradually degrade model performance and complicate the detection of malicious clients [39, 35], as the impact of attacks increases with the degree of non-IID [19, 26].
In this study, we introduce FedCC, a defense mechanism that mitigates untargeted model poisoning and targeted backdoor attacks in both IID and non-IID settings. FedCC leverages the Centered Kernel Alignment (CKA) of Penultimate Layer Representations (PLRs) to distinguish malicious clients from benign ones. PLRs are known to be highly specific to local data distributions [36], making clients distinguishable, especially when data is non-IID. Exploiting CKA’s well-known capability of measuring similarity in high-dimensional space, such as neural network representations, we accurately measure the similarity between the PLR of different clients, further allowing more accurate detection of malicious clients. Empirical evidence supporting our rationale is shown in Figure 2 and Table 1. The use of PLRs shows the highest separability between benign and malicious clients in all settings, as seen by the largest distance between clusters in the PLR. The use of CKA, specifically kernel CKA, is justified by its highest performance in both targeted and untargeted IID and non-IID settings. Furthermore, our method maintains data privacy by using similarity measures between local and global models, not having shared data. We named our proposed method FedCC, Federated learning using Clustered CKA.
@@
Empirical experiments show that FedCC outperforms existing defenses that rely on first-order statistics, which usually suffer from a high false negative detection rate. It significantly improves global accuracy against untargeted attacks (66.27% compared to similar approaches) and reduces attack confidence in targeted backdoor attacks to almost zero while preserving main task accuracy. These results highlight FedCC’s effectiveness and robustness in mitigating attacks in FL settings, particularly in non-IID data settings.
@@
  • We propose FedCC, a novel and scalable defense method against model poisoning attacks using the CKA of PLRs of local and global models and performing layer-wise weighted aggregation of their model parameters.
@@
  • We justify that PLR is a highly distinguishable feature to detect malicious clients by measuring the discrepancy between two clusters of clients.
@@
  • We demonstrate that CKA provides the most sensitive and accurate similarity measure between differently trained models when the server has no knowledge about clients or centralized data, even in non-IID settings.
@@
  • We empirically demonstrate the effectiveness of FedCC through exhaustive experiments. Our results show that comparing the CKA of PLRs filters malicious clients more reliably than the first-order statistics. FedCC outperforms existing defenses even in non-IID settings.
2 Backgrounds and Related Works
2.1 Poisoning Attacks in FL
@@
FL is vulnerable to poisoning attacks due to its distributed nature, where the server cannot directly inspect dataset quality or model integrity. Model poisoning attacks are particularly destructive, as adversaries can stealthily manipulate local model parameters to degrade global model performance [3, 5], and this paper focuses on such attacks.
In model poisoning attacks, client-side adversaries alter local model parameters before submitting them to the server. To avoid detection and prevent global model divergence, attackers optimize the local models for both training loss and an adversarial objective [5, 7]. For instance, A3FL [37] dynamically fine-tunes backdoor triggers to make them harder to detect, while IBA [24] generates robust triggers using a generative network, exploiting the global model as a discriminator. Attackers can further manipulate hyperparameters, such as learning rate, local epochs, batch size, and regularization [17, 21], dynamically before and during local training to evade detection. In a similar vein, DBA [33] embedded split triggers into local training data, associating them with targeted incorrect predictions. Unlike centralized backdoor attacks, DBAs distribute malicious updates across multiple adversarial participants, making them harder to detect with anomaly detection techniques.
Model poisoning attacks can be classified into untargeted and targeted types. Untargeted attacks degrade global model accuracy, while targeted attacks mislead the model to misclassify specific inputs without affecting other classes. Our defense mechanism mitigates both untargeted attacks by restoring global model accuracy and targeted backdoor attacks by overcoming the stealthiness and detection difficulty [31, 29].
2.2 Robust Aggregation algorithm in FL
@@
Several Byzantine-robust aggregation methods, based on summary statistics or anomaly detection, have been proposed to mitigate model poisoning and backdoor attacks in FL. Krum [4], for instance, selects the update with the smallest Euclidean distance to others but assumes IID data and overlooks outliers [2]. Multi-Krum, which averages updates from multiple clients, is known to be more effective in non-IID settings [4]. Similarly, Median [34] computes the coordinate-wise median, Trimmed Mean [34] excludes extreme values, and Bulyan [10] combines Krum and Trimmed Mean for added robustness. These methods, however, require knowledge of the number of attackers, often unavailable in practice.
Inspired by the observation that malicious clients’ model parameters exhibit higher similarity, Foolsgold [8] identified Sybils by measuring cosine similarity between gradients. While effective against multiple Sybils, it struggled when only a single malicious client exists or with IID data, where it can overfit. FLTrust [6] addressed this by assigning trust scores based on ReLU-clipped cosine similarities but risking privacy by raw data sharing. Similarly, Lockdown [11] improved robustness by pruning parameters unused by the majority of clients during training, while FLIP [38] rejected low-confidence samples at test time in addition to the adversarial training.
Despite these advances, many methods rely on summary statistics (mean or median), direction (cosine similarity), or distance (Euclidean) of weight vectors. However, they often struggle to differentiate benign clients with non-IID data from malicious ones, leading to misclassification and suboptimal performance. The limited performance might be partially attributed to the inadequacy of Euclidean geometry for comparing neural network representations.
In contrast, our approach uses Kernel CKA, a kernel-based metric, to measure the similarity between global and local models. This method effectively identifies compromised model parameters among clients with diverse data distributions, enhancing defense against model poisoning in FL and overcoming the shortcomings of previous techniques.
2.3 Penultimate Layer Representation in FL
@@
The penultimate layer is a neural network model’s second-to-last layer (i.e., before the softmax layer). Wang et al. [30] discovered that PLRs are effective in distinguishing malicious models from benign ones; within benign clients, benign PLRs follow the same distribution, whereas malicious PLRs do not, across datasets and neural network architectures. Specifically, they demonstrated that distances among benign PLRs are smaller than those of benign and malicious PLRs. Their proposed method, FLARE, assigned a trust score to each client based on pairwise PLR discrepancies defined by Maximum Mean Discrepancy among all model updates, allocating lower values to those farther from the benign distribution. The model updates are then scaled and averaged, weighted by each client’s trust score. Meanwhile, FLARE continuously redirects the global model using server-owned raw data, increasing the risk of data leakage and jeopardizing data privacy.
2.4 Centered Kernel Alignment
@@
Centered Kernel Alignment (CKA) [13] is a highly accurate similarity metric that measures how similar two differently initialized or trained neural networks are, producing a value between 0 (not similar) and 1 (identical). CKA can be calculated by: $CKA(X,Y)=\frac{HSIC(X,Y)}{\sqrt{HSIC(X,X)HSIC(Y,Y)}}=\frac{{\lVert K(X^{T})K(Y)\rVert}^{2}}{{\lVert K(X^{T})K(X)\rVert}{\lVert K(Y^{T})K(Y)\rVert}}$ where $HSIC$ is Hilbert-Schmidt Independence Criterion [9] (i.e., a not normalized version of CKA) and $K$ is the RBF kernel. It constructs similarity kernel matrices in the weight space and compares these matrices to characterize the representation space, allowing comparisons between layers with varying widths and different initialization schemes.
@@
CKA includes several vital properties for similarity measures of neural networks. It is non-invariant to invertible linear transformations, meaning similarity scores change if such transformations are applied. This property is vital; the gradient descent algorithm is not invariant to these transformations because similarity metrics invariant to linear transformations are inaccurate on models trained with gradient descent. It is also invariant to orthogonal transformations and isotropic scaling, which are desirable for neural networks trained with gradient descent owing to their stochastic nature. These properties make CKA a precise and reliable metric for measuring similarities between neural networks, particularly when comparing potential models from malicious clients.
@@
3 Threat Model
3.1 Attackers’ Goals
We consider two primary attack strategies in the context of FL: untargeted model poisoning and targeted backdoor attacks. Untargeted attacks (Fang attacks [7]) aimed to evade robust aggregation rules like Krum and Coordinate-wise Median or Trimmed Mean. Attackers manipulate the model parameters to degrade the global model’s overall accuracy. Targeted backdoor attacks, represented by Bhagoji et al. [3] and DBA [33], aim to deceive the global model into misclassifying specific data samples the attacker chose. The objective is to assign a different label chosen by the attacker while maintaining high accuracy for the remaining classes, making the attack more inconspicuous.
3.2 Attackers’ Capability
@@
  • An attacker has knowledge about compromised clients’ data, such as current local model, previous global model, and hyperparameters.
@@
3.3 Attack Strategy
3.3.1 Untargeted Model Poisoning Attacks [7]
@@
specifically aim to break Krum and Coordinate-wise Median (Coomed) aggregation rules, known to be byzantine failure tolerant. Their primary objective is to disrupt the model training process, thereby diminishing global test accuracy. The attacks involve manipulating model parameters, such as flipping the sign of malicious parameters, to steer the model in the opposite direction from its uninterrupted trajectory. Specifically, we denote the attack against Krum as Untargeted-Krum and the attack against Coomed as Untargeted-Med.
@@
Untargeted-Krum [7] alters malicious parameters resembling benign ones to maximize the chances of being selected by Krum. Specifically, the optimization problem is to find the maximum $\lambda,s.t.$ $w_{1}=Krum(w_{1},...,w_{m},w_{m+1},...,w_{n}),w_{1}=w_{G}-\lambda s,w_{i}=w_{1},$ for $i=2,3,...,m$ where $m$ is the number of attackers, $n$ is the total number of selected clients, $w_{G}$ is the previous global model, and $s$ is the sign of the global model parameter with no attack. The upper bound of the $\lambda$ can be solved as follows: $\lambda\leq\frac{1}{(n-2m-1)\sqrt{d}}\cdot\min_{m+1\leq i\leq n}\sum_{l\in\Gamma_{w_{i}}^{n-m-2}}D(w_{l},w_{i})+\frac{1}{\sqrt{d}}\cdot\max_{m+1\leq i\leq n}D(w_{i},w_{G})$ where $d$ is the number of parameters in the global model, $\Gamma$ is the set of $n-m-2$ benign local models having the smallest Euclidean distance to $w_{i}$, and $D$ is the Euclidean distance. $\lambda$ is halved until one of the compromised models is selected, or $\lambda$ is less than a threshold.
@@
Untargeted-Med [7] manipulates the model parameters based on the maximum and minimum so that chosen coordinate-wise median values direct toward an inverse direction. The attack starts with defining $w_{max,j}=\text{max}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$, which indicates the maximum, and $w_{min,j}=\text{min}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$ which indicates the minimum of the $j$th local model parameters on the benign clients, respectively. Also, to avoid sampled $m$ numbers being outliers, if $s_{j}=-1$, $m$ numbers in [$w_{max,j},b\cdot w_{max,j}$] (when $w_{max,j}>0$) or [$w_{max,j},b/w_{max,j}$] (when $w_{max,j}\leq 0$) are randomly sampled. Otherwise, $m$ numbers [$w_{min,j}/b,w_{min,j}$] (when $w_{min,j}>0$) or [$b\cdot w_{min,j},w_{min,j}$] (when $w_{min,j}\leq 0$) are randomly sampled. We set $b=2$ as the same as the paper [7].
3.3.2 Targeted Backdoor Attacks